OpenStudy (anonymous):

Find the geometric mean of the pair of numbers. 3 and 6

OpenStudy (anonymous):

same as last one' multiply them , take the square root

OpenStudy (anonymous):

\(c^2=a\times b\) c is the geometrical mean of a and b

OpenStudy (anonymous):

only this time you do not get a prefect square like the last one

OpenStudy (anonymous):

use that or this\[{a\over c}={c\over b}\]

OpenStudy (anonymous):

yeah I did and it came out as 18 then when I got 4.24 but I cant figure out what way to write that out

OpenStudy (anonymous):

\[\sqrt{9\times 3}=\sqrt{9}\times \sqrt{3}=3\sqrt{3}\]

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

or use a calculator if you want a decimal approximation

OpenStudy (anonymous):

i did use a calculator I am saying I dont know what way to write it out let me show you the options @satellite73

OpenStudy (anonymous):